If the 5th term of an arithmetic sequence is 32 and the 10th term is 57, find the following;
A) d
B) A1
C) Formula for An
D) A22

Answer by richwmiller(17219)

an = a1 + (n - 1)d
a5 = 32=a1 + 4d
a10=57 = a1 + 9d
57=a + 9d
32=a + 4d
25=5d
A) d=5
32=a1 + 4*5
B) a1=12
57=a1 + 9*5
57=a1+45
B) 12=a1
an = a1 + (n - 1)d
an = 12 + 5(n - 1)
an = 12 + 5n - 5
C) an = 5n +7
D) a22 = 5*22 +7=117